Please help. Use the bond term's below to answer the question
Maturity 6 years
Coupon Rate 3%
Face value $1,000
Annual Coupons
When you buy the bond the interest rate is 4%

Right after you buy the bond, the interest rate changes from 4.00% to 2.75% and remains there.

What is the price effect in year 5 ?
Answer #1

Price of bond when Interest rate is 4% = 30/(1.04) + 30/(1.042)+30/(1.043)+30/(1.044)+30/(1.04)5+ 1000+30/(1.046) = $947.58

Price of bond when Interest rate changes to 2.75% = 30/(1.0275) + 30/(1.02752)+30/(1.02753)+30/(1.02754)+30/(1.02755)+ 1000+30/(1.02756) = $1013.66

Hence price effect = $1013.66 - $947.58 = $66.08

Due to fall in Interest rate there will be rise in the price of bond which means lower return.Price and Interest both are inversly related to each other.